# references.bib

### v1

@@ 4.3 Robot Pose Optimization with Bayesian Optimization
Bayesian Optimization is a sampling-based, gradient-free optimization algorithm designed to find parameters $\mathbf{\mathbf{w}}^{*}$ that maximize a target function $f_{\text{BO}}(\mathbf{w})$. It achieves this by modeling $f_{\text{BO}}$ using a surrogate model—typically a Gaussian Process (GP) or random forest—that approximates the underlying function based on observed samples. At each iteration, BO selects new samples from $f_{\text{BO}}$ by maximizing an acquisition function, such as the Upper Confidence Bound (UCB) [srinivas2009gaussian], which balances exploration and exploitation. The acquisition function prioritizes regions of the parameter space that are both uncertain and potentially high-scoring, enabling BO to identify promising regions efficiently with a limited number of queries. To optimize the policy optimization score $K(p)$, we initialize BO with $N^{(\text{BO})}_{\text{init}}$ randomly sampled base poses in the scene and evaluate their corresponding scores. We then run $N^{(\text{BO})}_{\text{iter}}$ optimization rounds where each round proposes $N^{(\text{BO})}_{\text{batch}}$ new poses based on the acquisition function. The highest-scoring pose found among all iterations is selected as the final pose for executing the manipulation policy. Please check out the implementation details of our method in the supplementary materials.
@@ 5 Experiments
In experiments, we aim to investigate the following key questions: [Q1] Do non-policy-aware baselines fail to navigate to base poses that enable successful policy execution? [Q2] Do methods trained with large-scale, in-domain navigation data exhibit limited generalization to out-of-distribution room layouts? [Q3] Does our proposed method outperform these baselines and consistently identify suitable base poses for successful policy execution? [Q4] Are our mobilization feasibility metrics predictive of the actual performance across different methods? [Q5] Can our method be effectively deployed in real-world settings?
@@
5.1 Simulation Experiments
Metrics. In addition to reporting the success rate of the baselines (see Section 3.3) and our method, we also report Manipulation Policy Performance, which is the success rate of the original, non-mobile robot policy per task without mobilization. Note that the success rates of the base policies vary significantly across tasks. We report the success rate and standard deviation across all tasks.
Results. We present quantitative results for our method in Figure 6. LeLaN (non-policy-aware) and VLFM (non-policy-aware) perform poorly because they do not consider which robot poses enable a successful policy rollout (see Figure 7 for qualitative results). This validates [Q1]. BC w/ Nav (policy-aware) performs poorly in all tasks, showing that it is difficult to train a working mobile manipulation policy that generalizes to out-of-distribution room layouts, validating [Q2]. Our method outperforms all baselines in all tasks, achieving better performance in navigating to a pose suitable for executing the downstream policy, resulting in higher task success, answering [Q3].
Relation to mobilization feasibility metrics. Baselines that are not policy-aware perform worst in Turn on Faucet, Turn on Stove, and Turn on Microwave. These are also tasks that have the lowest spatial mobilization feasibility scores (Section 3.1). This suggests that poor spatial tolerance of a manipulation policy makes it challenging for a non-policy-aware method to perform well. Our method’s performance also correlates with the visual mobilization feasibility metric. It matches or exceeds Manipulation Policy (Perturb 5 cm) across tasks, except Close Drawer, which has low visual feasibility. As our method relies more on visual cues, this explains the performance drop, supporting [Q4].
@@ 5.1 Simulation Experiments
Ablations and analysis. We conduct ablation studies in two simulation environments, Close Door and Turn on Microwave. Specifically, we wish to understand which visual features are optimal when being used for selecting optimal robot base placement poses during policy mobilization. We consider the following ablated methods: (1) Ours w/o Feature Descriptor: this ablated method uses a DINO embedding vector instead of the DINO dense feature descriptor in $K_{\text{id}}$ to represent information in an image. (2) Ours w/ Policy Encoder: this method uses the ResNet-18 [he2016deep] encoder of the given policy to represent an image in $K_{\text{id}}$.
Figure 7: Visualization of qualitative performances in simulated experiments. This plot shows achieved navigation end poses of the robots produced by multiple competing methods in a sample scene in the Turn on Microwave task. The background plots the top-down map of the environment where the robot needs to run the designated policy. Each arrow represents a robot pose achieved by a method after navigation. The colors of the arrows represent method names. Semi-transparent arrows indicate failures and opaque arrows indicate successes. In the plot, our method consistently reaches the microwave door at an appropriate heading, while baselines fail to do so.
@@
We also design visualization tools to show the qualitative performance of competing methods. In Figure 7, we show top-down maps of the simulated benchmark environment overlaid with arrows to illustrate the navigation end poses achieved by different methods. See more visualizations in the supplementary materials.
@@ 5.2 Real Robot Experiments
To validate [Q5], we show a series of real robot experiments where a mobile robot performs various manipulation tasks in a mock grocery store. We pick three tasks that involve diverse objects and contact modes as pictured in Figure 9.
Figure 8: Real-robot experiment results. For each task and method evaluated, we track the incidence rates of certain performance milestones (i.e., navigation success, grasp success, full success). We also track the percentage of collision-free episodes. We consider a navigation attempt successful if the robot drives to a pose that not only faces the target object, but also is within a distance of 50 cm from the ideal target pose. A grasp success is defined to be when the robot stably grasps the object of interest (i.e. the chip bag, the paper towels, or the tray of tomatoes), and full success is counted only if the entire task is completed.
@@
Data collection and training. We collect 30 human-teleoperated demonstrations for the Pick Chips and Grab Paper Towels task and 50 for the Pour Tomatoes task. We use a Basler camera mounted on the mobile robot to collect visual information and record both the images and the base and arm poses at a frequency of 10 Hz. We use the collected data to train a Diffusion Policy [chi2023diffusionpolicy]. See supplementary materials for more robot setup and training details.
@@
Figure 9: Real-world tasks. (1) Pick Chips: the robot retrieves a deformable bag of chips from a shelf. (2) Bimanual Grab Paper Towels: the robot squeezes a pack of paper towels on a shelf between the two arms using non-prehensile manipulation, lifts it up, and pulls it off the shelf. (3) Pour Tomatoes: the robot picks up a tray of tomatoes on the left side of a produce bin, moves rightward via base motions, and pours the tomatoes from the tray into a basket. See videos on our website.
@@
Baselines and rollouts. We initialize the robot at 10 random poses located 0.5 to 1.5 meters from the target policy execution pose, with the constraint that the target scene must remain within the robot’s field of view. From each initialization, we execute both our proposed method and the baseline methods. For real-world experiments, we compare against one policy-aware and one non-policy-aware baseline. As the policy-aware baseline, we use BC w/ Nav. To construct a strong non-policy aware baseline, we introduce a Human baseline: participants are asked to manually drive the robot to the location they believe will maximize task success, without being given any information about how the manipulation policy was trained. Additional details about the human baseline are provided in the supplementary materials.
@@
Quantitative results. Quantitative results from the real-robot experiments are shown in Figure 8. The Human baseline displays an interesting pattern: although participants successfully navigate the robot near objects of interest, their lack of knowledge about the manipulation policy’s capabilities often results in suboptimal base poses, leading to low success rates. The BC w/ Nav baseline achieves higher performance but frequently positions the robot inappropriately–too close or poorly oriented realtive to the target object–resulting in failed manipulation attempts. In contrast, our method consistently outperforms both baselines, reliably identifying base poses that enable successful policy execution.
@@
Qualitative results. Aside from quantitative results, we also demonstrate real-world deployment of our method in a variety of interesting ways (see Figure 10).
@@
Finally, we show that our method can be easily extended to adapt to unseen object heights. To achieve this, we simply add one more delta robot height parameter to the optimization process, and command the robot to move to the target height instructed by our method before starting policy execution. We find that our method can flexibly determine the proper height to run the Bimanual Grab Paper Towel policy when the object is placed on shelves with different heights.
@@ 6 Conclusion
In this work, we presented Mobi-$\pi$, a framework for studying policy mobilization. We introduced a formal problem definition of policy mobilization, presented a spatial and visual metric for quantifying the feasibility of mobilization, and provided an evaluation platform consisting of simulation benchmarks, visualization tools, and baselines. We showcased a promising approach to solve this problem leveraging a 3D Gaussian Splatting model, a hybrid score function, and sampling-based pose optimization to identify suitable robot poses for policy execution. We demonstrated our method in simulated kitchen environments as well as a real-robot setup in a mock-grocery environment. This work provides a robust pathway for utilizing countless non-mobile robot policies on mobile platforms, enabling robots to perform intricate tasks in larger and more diverse environments.
@@ 7 Limitations and Discussions
Dynamic scene changes. Our approach assumes a static environment during policy execution and does not support dynamic scene updates. This limitation becomes apparent when chaining multiple policies in the same workspace: after executing one policy, changes in the environment are not reflected in the scene model, potentially leading to degraded performance in subsequent steps. Addressing this limitation would require integrating a dynamic neural rendering model that can be updated in real-time as the robot interacts with the scene [wu20244d, zheng2025gstargaussiansurfacetracking].
@@ Acknowledgments
We thank members of the Mobile Manipulation Team at Toyota Research Institute for building the robot used in this work and maintaining related infrastructures. We thank members of the Stanford IPRL lab, Leonidas J. Guibas, Sergey Levine, Shuran Song, Zi-ang Cao, Yufei Ding, Ray Song, Chen Wang, and Haoyu Xiong for their helpful discussions.
[…]
@@ A.1 Motivation and Alternative Approaches
Why should I care about policy mobilization? Why not simply train a mobile manipulation policy from a large dataset? Indeed, one approach to learning mobile manipulation is to train a policy from a dataset that includes both navigation and manipulation data. However, training such a policy typically requires large amounts of training data, since the policy needs to not only generalize to different navigation and manipulation scenarios but also seamlessly coordinate navigation and manipulation. We showed in our simulation experiments that the BC w/ Nav baseline fails to perform well in unseen room layouts despite learning from 5×5$\times$5 × more training data (see Section 5.1). Compared to training an end-to-end mobile manipulation policy, our policy mobilization framework provides a more data-efficient approach to learning mobile manipulation.
@@ A.2 Method
How sensitive are DINO dense descriptors to viewpoint changes? In the early stage of the project, we compared several feature representations including DINO dense descriptors, DINO features, policy encoder features, ResNet-50 [he2016deep], and DeepLab v3 [chen2017rethinking]. We empirically found that the DINO dense descriptor performs the best in finding an in-distribution image out of a set of similar-looking images rendered from various viewpoints. In the paper, we showed quantitative ablation results where our method outperforms variations of our method where the DINO dense descriptors are replaced with flat DINO features and policy encoder features. These observations coincide with findings in prior work [amir2021deep, hadjivelichkov2023one] that DINO dense descriptors carry rich, well-localized semantic and positional information. Also, note that our method is compatible with any improved visual feature representation that emerges in the future.
Figure 11: Sample 3D Gaussian Splatting renders. As seen in the samples, our 3DGS models synthesizes novel views that are imperfect. Nevertheless, our method maintains good performance.
@@
The proposed method builds a 3D Gaussian Splatting model. Does it scale to larger, more complex scenes? In simulation, we showed that our method effectively finds optimal base positions in kitchen scenes as large as $36m^{2}$. This proves that our method is useful in room-scale settings. To make our method useful in a multi-room setting, one can consider designing a system that first navigates to the correct room and then executes our method. In the real world, another bottleneck is building a high-quality 3D Gaussian Splatting model. Prior work [yu2024language] studied building 3DGS models for complex scenes by incrementally registering images captured using a multicamera system to minimize drift during scene mapping. Integrating methods like this into our framework is a clear direction for future work.

### v2

@@ 4 A Novel Method for Policy Mobilization
Bayesian Optimization is a sampling-based, gradient-free optimization algorithm designed to find parameters $\mathbf{\mathbf{w}}^{*}$ that maximize a target function $f_{\text{BO}}(\mathbf{w})$. It achieves this by modeling $f_{\text{BO}}$ using a surrogate model—typically a Gaussian Process (GP) or random forest—that approximates the underlying function based on observed samples. At each iteration, BO selects new samples from $f_{\text{BO}}$ by maximizing an acquisition function, such as the Upper Confidence Bound (UCB) [82], which balances exploration and exploitation. The acquisition function prioritizes regions of the parameter space that are both uncertain and potentially high-scoring, enabling BO to identify promising regions efficiently with a limited number of queries. To optimize the policy optimization score $K(p)$, we initialize BO with $N^{(\text{BO})}_{\text{init}}$ randomly sampled base poses in the scene and evaluate their corresponding scores. We then run $N^{(\text{BO})}_{\text{iter}}$ optimization rounds where each round proposes $N^{(\text{BO})}_{\text{batch}}$ new poses based on the acquisition function. The highest-scoring pose found among all iterations is selected as the final pose for executing the manipulation policy. See implementation details of our method in the supplementary materials.
5 The ramework
To study policy mobilization in more depth, we propose the ramework, which includes metrics to quantify the difficulty of mobilizing a given policy (Section 5.1), a suite of simulated mobile manipulation tasks to benchmark methods and policies for mobilization (Section 5.2), as well as several visualization tools.
5.1 Mobilization Feasibility Metrics
Given a policy for a task, we want to know how hard it is to mobilize it. We say a policy for a task is “difficult to mobilize” if running this policy on a mobile platform results in a large performance drop relative to its in-distribution performance. To quantify mobilization feasibility, we propose two distinct and complementary metrics as part of the ramework.
Visual mobilization feasibility metric. The visual appearance of the tasks can also affect mobilization difficulty. For example, if a policy manipulates an object while barely seeing the object in view, it will be harder to find a good location for policy execution. To quantify this, we evaluate how relevant objects of the task are represented in the scene. Formally, we sample $N_{v}$ images of the scene at near-optimal poses and compute the average percentage $S_{v}$ of the images occupied by the objects of interest in the task. The visual mobilization feasibility metric can be evaluated for methods that consider aligning the robot’s viewpoint with the policy’s capabilities during base placement. Naive baselines like LeLaN [17] and VLFM [18] are not policy-aware, so this metric does not apply.
5.2 Simulated Task Suite for Evaluating Policy Mobilization
Part of the ramework is a suite of simulated mobile manipulation tasks based on RoboCasa [20] that will allow us to study the policy mobilization problem and benchmark different methods. Our task suite includes five challenging tasks illustrated in Figure 3. For each task, we train a standard image-based imitation learning policy [32] from 300 demonstrations. These demos are generated using MimicGen [83] with procedurally generated textures in 5 training room layouts. In each task, we use a Franka Panda robot mounted on an Omron mobile base initialized from the same base pose. Each task comes with a language description and a horizon of 500 steps. The goal of the benchmark is to deploy these policies into 5 held-out scenes with unseen room layouts and textures. Note that since these tasks come originally from RoboCasa, the tasks and policies are not designed for the problem of policy mobilization. Therefore, this benchmark can serve as an unbiased challenge task suite for testing methods for policy mobilization.
Figure 3: A suite of simulated tasks for benchmarking performance of policy mobilization methods. We pick five single-stage manipulation tasks from the RoboCasa benchmark [20]: (1) Close Door: push or close the microwave or cabinet door using the robot gripper. (2) Turn on Faucet: precisely move the faucet handle to turn on water. (3) Turn on Microwave: push a small button on the microwave. (4) Close Drawer: close an open drawer next to the robot. (5) Turn on Stove: turn a specific knob out of six knobs on the stove. Here, we visualize two successful rollouts for each task. Note that episodes are executed in environments with different layouts and textures. In the Close Door task, the robot might encounter either a cabinet door or a microwave door. The Turn on Faucet, Turn on Microwave, and Turn on Stove tasks require accurate robot placement to be executed correctly. In the Close Drawer task, the drawer can be either on the left or right side of the robot.
Quantifying mobilization feasibility of benchmark tasks. We plot the spatial mobilization feasibility metrics of the policies in the five simulated kitchen tasks in Figure 4. We observe that the most challenging task to mobilize based on our metric is the Turn on Stove task ($\phi=0.031$). This reflects the difficulty to precisely turn a small knob which has limited contact area. Figure 5 shows the visual mobilization feasibility metric evaluated in the five environments. Results show that visual mobilization feasibilities do not necessarily match spatial mobilization feasibilities. More specifically, the Close Drawer task, which appears to be easy to mobilize spatially, turns out to be the most difficult in terms of the visual mobilization feasibility metric. This is expected because even in spatially tolerant tasks, using proper visual features to find optimal policy initialization poses can be difficult. In experiments, we show how the mobilization metrics are correlated with the method performance.
@@ 6 Experiments
We aim to investigate how our policy-aware method performs compared to other policy-aware and non-policy-aware baselines, if our mobilization feasibility metrics are predictive of method performance, and if our method can be reliably deployed in real-world settings.
Baselines. We divide baselines into two categories. The first type of baselines navigates to the object of interest without considering the manipulation policy’s capabilities. These methods are not policy-aware. Representative methods include: (1) LeLaN [17], a language-conditioned navigation model trained from large-scale videos. Since the pre-trained LeLaN model is trained on only real-world data, we fine-tune the pre-trained LeLaN model with 7.5k episodes of in-domain, randomized navigation data in simulation. At test time, we run LeLaN for a fixed number of steps to approach the object while using the scene point cloud for collision checking, and then execute the target policy. (2) VLFM [18], a zero-shot navigation model that explores the scene, uses a ground-truth detector to locate the object, selects a point on its unprojected point cloud, and navigates to it using RRT and the scene point cloud for collision checking. As shown in our experiments, these types of methods often fail to find suitable base poses as they are not taking manipulation policy performance into account, resulting in failed policy executions.
The second type of baselines is policy-aware and leverages large-scale data to connect navigation with manipulation. A representative method is BC w/ Nav: a Behavior Transformer [32] trained to jointly perform navigation and manipulation using combined demonstrations. We provide 1,500 demonstrations per task, $5\times$ the number of demos used compared to our base manipulation policy. We show in experiments that these methods generalize poorly to unseen layouts despite extensive data collection. See supplementary materials for baseline details.
6.1 Simulation Experiments
Metrics. In addition to reporting the success rate of the baselines and our method, we also report Manipulation Policy Performance, which is the success rate of the original, non-mobile robot policy per task without mobilization. Note that the success rates of the base policies vary significantly across tasks. We report the success rate and standard deviation across all tasks.
@@ 6.1 Simulation Experiments
Results. We present quantitative results for our method in Figure 6. LeLaN (non-policy-aware) and VLFM (non-policy-aware) perform poorly because they do not consider which robot poses enable a successful policy rollout (see Figure 7 for qualitative results). BC w/ Nav (policy-aware) performs poorly in all tasks, showing that it is difficult to train a working mobile manipulation policy that generalizes to out-of-distribution room layouts. Our method outperforms all baselines in all tasks, achieving better performance in navigating to a pose suitable for executing the downstream policy, resulting in higher task success.
Relation to mobilization feasibility metrics. Baselines that are not policy-aware perform worst in Turn on Faucet, Turn on Stove, and Turn on Microwave. These are also tasks that have the lowest spatial mobilization feasibility scores (Section 5.1). This suggests that poor spatial tolerance of a manipulation policy makes it challenging for a non-policy-aware method to perform well. Our method’s performance also correlates with the visual mobilization feasibility metric. It matches or exceeds Manipulation Policy (Perturb 5 cm) across tasks, except Close Drawer, which has low visual feasibility. As our method relies on visual cues provided by the target objects, this explains the performance drop.
@@
Ablations and analysis. We conduct ablation studies in two simulation environments, Close Door and Turn on Microwave. We consider the following ablated methods: (1) Ours w/o Feature Descriptor: this ablated method uses a DINO embedding vector instead of the DINO dense feature descriptor in $K_{\text{id}}$ to represent information in an image. (2) Ours w/ Policy Encoder: this method uses the ResNet-18 [84] encoder of the given policy to represent an image in $K_{\text{id}}$.
Figure 7: Visualization of qualitative performances in simulated experiments. This plot shows achieved navigation end poses of the robots produced by multiple methods in a sample scene in the Turn on Microwave task. The background plots the top-down map of the scene where the robot needs to run the designated policy. Each arrow shows a robot pose achieved by a method after navigation. The colors of the arrows represent method names. Semi-transparent arrows indicate failures and opaque arrows indicate successes. In the plot, our method consistently reaches the microwave door with the appropriate heading, while baselines fail to do so.
@@
We also design visualization tools to show the qualitative performance of competing methods. In Figure 7, we show top-down maps of the simulated benchmark environment overlaid with arrows to illustrate the navigation end poses achieved by different methods. See more ablations and analysis in the supplementary materials.
@@ 6.2 Real Robot Experiments
To validate that our method can be effectively deployed in real-world settings, we show a series of real robot experiments where a mobile robot performs various manipulation tasks in a mock grocery store (see Figure 8).
@@
Data collection and training. We collect 30 human-teleoperated demonstrations for the Pick Chips and Grab Paper Towels task and 50 for the Pour Tomatoes task. We use a Basler camera mounted on the mobile robot to collect visual information and record both the images and the base and arm poses at a frequency of 10 Hz. We use the collected data to train a Diffusion Policy [1]. See supplementary materials for more robot setup and training details.
@@
Figure 8: Real-world tasks. We pick three tasks that involve diverse objects and contact modes. (1) Pick Chips: the robot retrieves a deformable bag of chips from a shelf. (2) Bimanual Grab Paper Towels: the robot squeezes a pack of paper towels on a shelf between the two arms using non-prehensile manipulation, lifts it up, and pulls it off the shelf. (3) Pour Tomatoes: the robot picks up a tray of tomatoes on the left side of a produce bin, moves rightward via base motions, and pours the tomatoes from the tray into a basket. See videos on our website.
@@
Baselines and rollouts. We initialize the robot at 10 random poses located 0.5 to 1.5 meters from the target policy execution pose, with the constraint that the target scene must remain within the robot’s field of view. From each initialization, we execute both our proposed method and the baseline methods. For real-world experiments, we compare against one policy-aware and one non-policy-aware baseline. As the policy-aware baseline, we use BC w/ Nav. To construct a strong non-policy-aware baseline, we introduce a Human baseline: participants are asked to manually drive the robot to the location they believe will maximize task success, without being given any information about how the manipulation policy was trained. See details about this baseline in the supplementary materials.
Figure 9: Real-robot experiment results. For each task and method evaluated, we track the incidence rates of certain performance milestones (i.e., navigation success, grasp success, full success). We also track the percentage of collision-free episodes. We consider a navigation attempt successful if the robot drives to a pose that not only faces the target object, but also is within a distance of 50 cm from the ideal target pose. A grasp success is defined as when the robot stably grasps the object of interest (i.e. the chip bag, the paper towels, or the tray of tomatoes), and full success is counted only if the entire task is completed.
@@
Quantitative results. Quantitative results from the real-robot experiments are shown in Figure 9. The Human baseline displays an interesting pattern: although participants successfully navigate the robot near objects of interest, their lack of knowledge about the manipulation policy’s capabilities often results in suboptimal base poses, leading to low success rates. The BC w/ Nav baseline achieves higher performance but frequently positions the robot inappropriately–too close or poorly oriented realtive to the target object–resulting in failed manipulation attempts. In contrast, our method outperforms both baselines, reliably identifying base poses that enable successful policy execution.
Qualitative results.
@@
Aside from quantitative results, we also demonstrate real-world deployment of our method in a variety of interesting ways (see Figure 10).
@@
Finally, we show that our method can be easily extended to adapt to unseen object heights. To achieve this, we simply add one more delta robot height parameter to the optimization process and command the robot to move to the target height instructed by our method before starting policy execution. We find that our method can flexibly determine the proper height to run the Bimanual Grab Paper Towel policy when the object is placed on shelves with different heights.
@@ 7 Conclusion
We introduced a framework for studying policy mobilization. We formalized the policy mobilization problem, proposed a new method for solving this problem, and introduced an evaluation suite with simulation benchmarks, visualizations, and baselines. We demonstrated our method in simulated kitchen environments as well as a real-robot setup in a mock-grocery environment. This work provides a robust pathway for utilizing countless non-mobile robot policies on mobile platforms, enabling robots to perform intricate tasks in larger and more diverse environments.
@@ 8 Limitations and Discussions
Dynamic scene changes. Our approach assumes a static environment during policy execution and does not support dynamic scene updates. This limitation becomes apparent when chaining multiple policies in the same workspace: after executing one policy, changes in the environment are not reflected in the scene model, potentially leading to degraded performance in subsequent steps. Addressing this limitation would require integrating a dynamic neural rendering model that can be updated in real-time as the robot interacts with the scene [85, 86].
@@ Acknowledgments
We thank members of the Mobile Manipulation Team at Toyota Research Institute for building the robot used in this work and maintaining related infrastructures. We thank members of the Stanford IPRL lab, Leonidas J. Guibas, Sergey Levine, Shuran Song, Zi-ang Cao, Yufei Ding, Ray Song, Chen Wang, and Haoyu Xiong for their helpful discussions. We thank Cherie Ho, Marion Lepert, and Kevin Lin for their help during the rebuttal.
@@ A.1 Motivation and Alternative Approaches
Why should I care about policy mobilization? Why not simply train a mobile manipulation policy from a large dataset? Indeed, one approach to learning mobile manipulation is to train a policy from a dataset that includes both navigation and manipulation data. However, training such a policy typically requires large amounts of training data, since the policy needs to not only generalize to different navigation and manipulation scenarios but also seamlessly coordinate navigation and manipulation. We showed in our simulation experiments that the BC w/ Nav baseline fails to perform well in unseen room layouts despite learning from $5\times$ more training data (see Section 6.1). Compared to training an end-to-end mobile manipulation policy, our policy mobilization framework provides a more data-efficient approach to learning mobile manipulation.
@@
Why not consider an image-goal navigation method or further achieve fine-grained navigation [87]? Image-goal navigation won’t easily apply because it is unclear what goal image to use. For example, in sim, the policy is deployed in novel environments with out-of-distribution layouts and textures, and it is trained on data from multiple diverse scenes. This means that the test-time scene does not match any scene captured during policy training. A similar argument applies to fine-grained navigation: one needs to first search for where to navigate to before running navigation.
@@ A.2 Method
How sensitive are DINO dense descriptors to viewpoint changes? In the early stage of the project, we compared several feature representations, including DINO dense descriptors, DINO features, policy encoder features, ResNet-50 [84], and DeepLab v3 [88]. We empirically found that the DINO dense descriptor performs the best in finding an in-distribution image out of a set of similar-looking images rendered from various viewpoints. In the paper, we showed quantitative ablation results where our method outperforms variations of our method where the DINO dense descriptors are replaced with flat DINO features and policy encoder features. These observations coincide with findings in prior work [89, 90] that DINO dense descriptors carry rich, well-localized semantic and positional information. Also, note that our method is compatible with any improved visual feature representation that emerges in the future.
Figure 11: Sample 3D Gaussian Splatting renders. As seen in the samples, our 3DGS models synthesize novel views that are imperfect. Nevertheless, our method maintains good performance.
@@
The proposed method builds a 3D Gaussian Splatting model. Does it scale to larger, more complex scenes? In simulation, we showed that our method effectively finds optimal base positions in kitchen scenes as large as $36m^{2}$. This proves that our method is useful in room-scale settings. To make our method useful in a multi-room setting, one can consider designing a system that first navigates to the correct room and then executes our method. In the real world, another bottleneck is building a high-quality 3D Gaussian Splatting model. Prior work [91] studied building 3DGS models for complex scenes by incrementally registering images captured using a multicamera system to minimize drift during scene mapping. Integrating methods like this into our framework is a clear direction for future work.
@@
Why does the proposed method use L2 distance in latent space as the in-distribution score rather than a distribution score (e.g. compute the exact likelihoods such as normalizing flows (NFs) as in [92])? While we could indeed search among more frames in the training dataset, doing so is computationally more expensive. We empirically found that only using the initial frames led to good performance. Using the L2 distance on DINO features is a simple, off-the-shelf, and empirically effective proxy for distribution matching that does not require training complex density models like normalizing flows (NFs), which require additional training, careful tuning, and might overfit on small datasets. We also note that the task suite we developed in sim can be used to develop alternative score functions that utilize more than just initial frames or employ flow-based methods.